the radius of gyration of a solid sphere of radius r about a certain axis is r. the distance of this axis from the centre of the sphere is √0.6r explain???

Grade:11

1 Answers

Unknown aspirant
5 years ago
Using parallel axis theorem ,
I =Icm(about the centre of mass) + M(mass of the sphere)h2.    where h is the required distance of the axis from centre of sphere
MK2 = 2/5 MR2. + Mh
K2= 2R2/5 + h2
R= 2R2/5. +  h2.       (given that.  K=R. )
3R2/5 = h2
Which gives h=√3/5R2
h=√0.6. R.   (Answer)
R
 
MK